OpenStudy (anonymous):

Which exponential function goes through the points (1, 32) and (4, 2048)? f(x) = 4(4)x f(x) = 8(2)-x f(x) = 8(4)x f(x) = 4(4)-x

OpenStudy (anonymous):

check which ones works for example \[f(x)=4\times 4^x\] does not work becaues \[f(1)=4\times 4^2=16\] not \(32\) as required

OpenStudy (anonymous):

so no, not A

OpenStudy (anonymous):

the answer is a